Week of 161219

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes




  • local: Ignacio (databases), Julia (WLCG), Maarten (SCOD + ALICE), Vincent (security)
  • remote: Andrew (NLT1), Dario (ATLAS), David (FNAL), Dimitri (KIT), Gareth (RAL), Luca (CNAF), Matteo (CNAF), Rolf (IN2P3), Sang-Un (KISTI), Stefano (CMS), Xin (BNL)

Experiments round table:

  • ATLAS reports (raw view) -
    • Production
      • Stable production levels using 250-300k cores
      • Regular data transfers after last week's problems with proxies in FTS
      • Continuing validation tests in view of big reprocessing
    • Problems
      • Almost all production and analysis jobs at INFN-T1 fail since Friday morning as they cannot see the input files - under investigation. Looks like a problem with an update to our pilots last week.
        • Luca: as a temporary workaround lcg-cp was used instead
      • Derivation productions using fast merge produce unreadable output and downstream jobs cannot open the files - being checked.
    • Plan
      • Fill the Grid for the holiday period mostly with simulation and keep an eye or two (best effort) on it.
    • Extra
      • Best wishes to everybody!

  • CMS reports (raw view) -
    • CMSINTR and CMSONR Oracle DB's switched over from P5 to CC on Wed. Dec 14/15: all OK
    • Stefano: plans for grid activities during the break are being finalized;
      many thanks to the sites and best wishes from CMS!

  • ALICE -
    • Expectations for the end-of-year break:
      • steady MC production
      • p-A raw data reconstruction
      • high analysis activity: preparations for Quark Matter conference, Feb 6-11
    • Thanks to all sites and experts for another successful year!
    • Season's greetings and best wishes for 2017!

  • LHCb reports (raw view) -
    • Activity
      • Data reconstruction (proton-ion), MC and user analysis.
    • Site Issues
      • T0:
        • NTR
      • T1:
        • SARA: Aborted pilots: (GGUS:125622). It is fixed now.
    • Merry Christmas and Happy new year!

Sites / Services round table:

  • ASGC: Scheduled downtime on 27 Dec. for upgrade of networking hardware and several services (https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=21977)
  • BNL: NTR
  • EGI:
  • FNAL:
    • FTS, PhEDEx and Xrootd redirector have all been upgraded
  • GridPP:
  • IN2P3: Late afternoon on last Monday at IN2P3-CC, an internal problem with the power supply of two racks led to a crash of about 400 jobs, mainly multicore ones of ATLAS, sorry for that.
  • JINR:
  • KIT:
    • we would like to have a short (~1h) downtime to fix a memory issue affecting the dCache SRM for CMS
      • the matter is being followed up via a ticket for CMS (GGUS:125634)
  • NDGF:
  • NL-T1:
    • SARA’s downtime last Thursday was unsuccessful; the Qfabric software update had to be rolled back. After the rollback, we had one rack of dCache pool nodes without connectivity. Because of this we extended the downtime. On Friday morning, the fibers of the uplink were cleaned and that solved the problem. It turned out that also other racks had dirty fibers on their uplinks. A few of them were cleaned on Friday; others will be cleaned today. As these uplinks are redundant, this operation should be transparent. We hope that cleaning the fibers will solve the network performance issue we’ve seen since the datacenter move.
    • SARA's batch scheduler had crashed this weekend; restart this morning. GGUS:125622
  • NRC-KI:
  • OSG:
  • PIC:
  • RAL: We will be updating Castor to version 2.1.15 in January. Dates are announced via GOC DB.

  • CERN computing services:
  • CERN storage services:
  • CERN databases:
    • the Oracle DB switches went OK for CMS and ALICE
    • we also intend to test that for ATLAS at some point
    • LHCb do not have a local replica
  • GGUS: Reminder! Year-End info on GGUS: During the Year End holiday period, if you observe a problem with GGUS and you cannot submit a ticket via https://ggus.eu, please contact the KIT site operator (info in https://goc.egi.eu/portal). The KIT Emergency telephone number is +49 721 60828383
  • Monitoring:
  • MW Officer: NTR
  • Networks:
  • Security: Security incident detected in EGI, EGI-20161214-01: see last week's broadcast (more to come later)


  • Notice: there will be a leap second on Dec 31 at 23:59:60 UTC
    • Past occurrences have triggered problems in various SW (OTG:0034791)

Season's Greetings!

  • THANKS for your help in making 2016 a very successful year for WLCG!
    • Further challenges and opportunities await us in 2017... smile

  • Next meeting: Mon Jan 9
Edit | Attach | Watch | Print version | History: r18 < r17 < r16 < r15 < r14 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r18 - 2016-12-19 - MaartenLitmaath
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback